OpenStudy (anonymous):

use the quadratic formula = 0.5x^2 = 2x + 0.5

OpenStudy (zzr0ck3r):

\[\frac{1}{2}x^2=2x+\frac{1}{2}?\]

OpenStudy (zzr0ck3r):

@fareastmovement13 ?